state true or false-if for a given current,resistance is changed from R to R/n,power consumed changes from P to nP
1
Expert's answer
2012-06-01T08:30:48-0400
P = V · I = (I · R) · I = I^2 · R
P0 = I^2 · R
P1 = I^2 · R /
n

Answer: false, power consumed changes from P to P/n.